# From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context

### v1

@@ 4.2 Prompt Construction
Given a node $u$, the projector outputs a matrix $\Pi(f_{\Phi}(u))=Z_{u}\in\mathbb{R}^{k\times h}$, which we refer to as a soft prompt embedding. This matrix is injected directly into the LLM’s embedding space and interleaved with natural language token embeddings.
For a node $v$ that we wish to explain, we consider its computation tree for GNN $\Phi$, denoted by $\mathcal{T}^{\phi}_{v}$. $\mathcal{T}^{\phi}_{v}$ is a tree with root node $v$ and depth $L$, the number of layers of $\Phi$, composed of all paths of length $L$ starting at $v$, with repetitions, concatenated to the root $v$. We construct the LLM input as a sequence of embeddings consisting of:
  • Standard token embeddings for the system prompt and user instruction.
  • A segment labeled “Target Paper Embedding Representation,” where we insert the soft prompt matrix $Z_{v}$ for the target node, bracketed by textual markers.
  • An enumeration of the nodes in $\mathcal{T}^{\phi}_{v}$, where each node is encoded by interleaved natural language markers and its own soft-prompt embeddings.
  • Final instructions text prompting the model to provide a summary and binary support decision for each computation tree node based on the embeddings.
The LLM sees the embedded GNN features as native tokens, enabling reasoning over both GNN’s representation and natural language inputs. We give an example of the resulting prompt in D.
@@ 4.3 Explanation Generation
To produce the explanations, the LLM first predicts whether each node in the computation tree supports or does not support the target node’s classification. This partitions the nodes of the computation tree into three categories: (i) Supporting nodes, (ii) Opposing nodes, (iii) Neutral nodes, not assigned to either category. The explanation subgraph is then formed by including all supporting nodes and a controlled number of neutral nodes. The inclusion of hallucinated nodes is proportional to the number of supporting nodes relative to the total size of the computation tree, as described in our size constraint formulation.
For each node $v$, the LLM generates a natural language explanation $E_{v}$ considering the influence of nodes of $\mathcal{T}^{\Phi}_{v}$ with regards to the classification of node $v$. We define a label attribution function based on the parsing of $E_{v}$:
$\chi:\mathcal{T}^{\phi}_{v}\to\{-1,0,1\}$
such that for each node $u\in\mathcal{T}^{\phi}_{v}$:
@@
To address uncertainty and enforce conciseness, we introduce a refinement function $\rho:2^{\mathcal{T}^{\phi}_{v}}\to 2^{\mathcal{T}^{\phi}_{v}}$ such that the final explanation subgraph is:
$S_{v}=\rho(S^{+}_{v},S^{-}_{v},S^{0}_{v})$
@@
Mitigating Hallucination. LLM-generated explanations may hallucinate, especially when multiple node embeddings are projected within a single prompt. We address this issue through two complementary mechanisms within our framework:
  • Prompt templating with constrained node sets: To reduce hallucinations, we ensure that our prompts contain a fixed set of nodes. We explicitly instruct the LLM to generate explanations only referencing these nodes, limiting the possibility of hallucinating unsupported nodes.
  • Post hoc filtering: After the LLM generates the explanations, any hallucinated nodes are filtered out in a post-processing step. This filtering process ensures that the final explanation remains relevant and grounded in the actual GNN model’s predictions, as detailed in Section 4.3.
@@
Cora WikiCS Liar Amazon
Fidelity Size Fidelity Size Fidelity Size Fidelity Size
Node 73.6% 1 77.8% 1 95.0% 1 73.8% 1
Random (50%) 93.5% 18.1 88.9% 19.5 95.0% 19.0 84.4% 7.24
Random (25%) 79.3% 10.0 86.9% 10.0 95.0% 18.13 73.3% 4.03
GNNExplainer 92.0% 18.38 95.4% 9.7 100.0% 12.4 95.5% 5.63
PGExplainer 86.5% 16.2 86.2% 21.2 95.0%* 1.00* 89.0% 5.73
TAGExplainer 60.0% 1.30 80.0% 1.20 95.0%* 1.00* 73.8% 1.52
LLM 93.3% 20.4 91.2% 9.8 95.0% 11.1 86.5% 5.07
Our method 94.6% 17.4 89.0% 8.9 94.8% 10.3 87.0% 1.30
@@ 5 Experiments
Our method, Logic, acts as a light-weight framework to generate explanations of Gnns especially for TAGs. We demonstrate the effectiveness of Logic in various settings. Our code is publicly available at: https://github.com/pbaghershahi/LOGIC.git.
5.1 Evaluation metrics
We use the following metrics in the experiments:
1) Fidelity: Fidelity captures how well an explainable model reproduces or approximates the GNN model logic through its prediction.
Mathematically, the fidelity of the explainer $\Psi$ of GNN $\Phi$ on dataset $\mathcal{G}$ over $\mathcal{V}$ nodes is:
$Fid=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in V}\mathds{1}(\Phi(v,\Psi(v,G))=\Phi(v,G))$
2) Explanation size: The explanation size measures the complexity or compactness of the explanation subgraph produced by the explainer $\Psi$. Specifically, it captures the number of nodes in the explanation for a given prediction. Smaller explanation sizes are generally preferred, as they are easier to interpret. The explanation size is:
$Size=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in\mathcal{V}}|\Psi(v,G)|$
@@ 5.2 Experimental setup
Datasets. We use real-world TAG datasets to evaluate the performance of our method: Cora Sen et al. (2008) - a co-citation network of research papers, WikiCS Mernyei and Cangea (2022) - a network of related Wikipedia articles, the Amazon-Product Feng et al. (2024) dataset - an E-commerce product co-purchase network, and the Liar Wang (2017) dataset - a fake news detection knowledge-graph. Additional details about the processing and relevance of the datasets are provided in Appendix B.
Models. We consider a GCN Kipf and Welling (2017a) as the base GNN model for the node classification task for all datasets. We train a model consisting of three convolutional layers, with the last one being the classification layer. The training/validation/testing split is $60\%/20\%/20\%$ for all the datasets. We give additional information about the GCN training in Appendix C. We use Meta’s Llama-3.1-8B-Instruct from the HuggingFace library as the frozen LLM in our experiments. We run all experiments on NVIDIA-GeForce-RTX-3090 GPUs, 24GB RAM.
Figure 2: A visualization of the explanations on the Amazon dataset. Left shows the explanations subgraph generated by GNNExplainer. Right shows the explanation generated by Logic. The node marked by a bold circle (product #7) is the node of interest for the explanation. Nodes with the same color are from the same category. Although both models have similar quantitative performance (in terms of size and fidelity), the LLM-generated explanations are textually clarified, making them more human-understandable.
Baselines. To compare our method’s performance with existing approaches, we consider the following five baselines: (1) Node: This baseline uses only the target node itself as the explanation subgraph. This represents the minimal possible context and serves as a lower bound on explanation size. (2) Random: We introduce a randomized baseline that selects a subgraph around each target node as its explanation subgraph. Specifically, for each node, we sample a subgraph of half the size of its computation tree. The results are averaged over 5 samples. (3-5) Graph Explainers: We compare Logic with GNNExplainer Ying et al. (2019a), PGExplainer Luo et al. (2020), subgraph-based explainers for node classification tasks, which do not involve any LLM-based component.
We add the recent LLM-based graph explanation method, TAGExplainer Pan et al. (2024). To the best of our knowledge, TAGExplainer is the only existing LLM-based explanation approach specifically designed for TAGs.
5.3 Quantitative Retults
Table 1 presents the results. One major observation is that on all the datasets, the Node baseline (target node as the singleton explanation subgraph) is considerably powerful, denoting how informative the textual attributes of each node are to guide the GNN’s predictions. Assessing based on fidelity, GNNExplainer, achieves the best performance over most of the datasets, except for Cora, on which Logic achieves better results. However, as discussed before, fidelity alone does not reflect the superiority of a method. Essentially, conventional GNNs (GCN in our case) score a sample, i.e., a node, based on the combination of its own information and the aggregated information of its neighbors, meaning that the context surrounding a node highly impacts its prediction score. Therefore, a GNN explainer that explains a target node with its whole induced subgraph achieves 100% fidelity. Meanwhile, the majority of nodes in the same neighborhood of graphs with a high homophily ratio belong to the same category. Thus, selecting a comparably large number of neighbors of a target node on these types of graphs can most probably guarantee the same representation for the target node as if the whole induced subgraph is selected.
Logic achieves a strong balance between fidelity and explanation size across all four datasets. Notably, Logic matches or exceeds the fidelity of traditional explainers like GNNExplainer, outperforming it on Cora, while producing substantially smaller explanation subgraphs, particularly on Liar and Amazon. This result is especially significant for TAGs, where textual node attributes carry rich semantic information; Logic effectively distills these semantics into compact, interpretable explanations without compromising on alignment with the original GNN predictions. In contrast, baseline methods such as Random or Node either inflate the explanation size or overly restrict the context. These results validate our design choice of using LLMs guided by projected GNN embeddings to construct context-aware yet concise explanation subgraphs, offering a superior trade-off between interpretability and predictive fidelity.
We present additional experiments for different GNN architectures and LLMs in Appendix E.
@@ 5.3 Quantitative Retults
Table 2: Expert qualitative evaluation of the explanations from Logic and GNNexplainers on Cora. $M_{1}$ denotes the mean scores for GNNexplainer and $M_{2}$ the mean scores for Logic.
Dimension $M_{1}\uparrow$ $M_{2}\uparrow$ $\Delta(M_{2},M_{1}$)
Understandability $2.9_{\scriptstyle\pm 0.8}$ $3.2_{\scriptstyle\pm 0.8}$ $0.3_{\scriptstyle\pm 0.8}$
Trustworthiness $2.7_{\scriptstyle\pm 0.8}$ $3.1_{\scriptstyle\pm 0.8}$ $0.4_{\scriptstyle\pm 0.8}$
Insightfulness $3.3_{\scriptstyle\pm 1.1}$ $3.1_{\scriptstyle\pm 1.1}$ $0.2_{\scriptstyle\pm 1.1}$
Satisfaction $2.2_{\scriptstyle\pm 0.9}$ $2.7_{\scriptstyle\pm 0.9}$ $0.5_{\scriptstyle\pm 0.9}$
Confidence $0_{\scriptstyle\pm 0.9} 3.3$$3_{\scriptstyle\pm 0.9}$ $0.3_{\scriptstyle\pm 0.9}$
Convincingness $3.3_{\scriptstyle\pm 1.2}$ $3.7_{\scriptstyle\pm 1.2}$ $0.4_{\scriptstyle\pm 1.2}$
Communicability $2.9_{\scriptstyle\pm 0.8}$ $3.2_{\scriptstyle\pm 0.8}$ $0.3_{\scriptstyle\pm 0.8}$
Usability $3.3_{\scriptstyle\pm 0.4}$ $3.5_{\scriptstyle\pm 0.4}$ $0.2_{\scriptstyle\pm 0.4}$

### v3

@@ 4.2 Prompt Construction
Given a node $u$, the projector outputs a matrix $\Pi(f_{\Phi}(u))=Z_{u}\in\mathbb{R}^{k\times h}$, which we refer to as a soft prompt embedding. This matrix is directly injected into the LLM’s embedding space and interleaved with natural language token embeddings.
For a node $v$ that we wish to explain, we consider its computation tree for GNN $\Phi$, denoted by $\mathcal{T}^{\phi}_{v}$. $\mathcal{T}^{\phi}_{v}$ is a tree with root node $v$ and depth $L$ (the number of layers of $\Phi$) composed of all paths of length $L$ starting at $v$, concatenated to the root $v$. We construct the LLM input as a sequence of embeddings consisting of: (i) Standard token embeddings for the system prompt and user instruction; (ii) The target node’s $k$-token embedding $Z_{v}$—generated after enclosing the node’s text descriptor by text markers; (iii) An enumeration of the nodes in $\mathcal{T}^{\phi}_{v}$, where each node is enclosed by text markers and its own soft-prompt embeddings; and (iv) Final query instructions. By treating the GNN embeddings as native tokens, the LLM reasons across both GNN representations and textual features (see Figure 2 and Appendix D.2).
@@ 4.3 Explanation Generation
First, the LLM predicts whether each node in the computation tree supports or does not support the target node’s classification. This partitions the nodes of the computation tree into three categories: (i) Supporting nodes, (ii) Opposing nodes, and (iii) Neutral nodes, which do not appear in either category. Formally, the LLM generates a natural language explanation $E_{v}$ based on $\mathcal{T}^{\Phi}_{v}$ and the classification of $v$. Given the parsing of $E_{v}$ we apply a label attribution function $\chi:\mathcal{T}^{\phi}_{v}\to\{-1,0,1\}$ for each node $u\in\mathcal{T}^{\phi}_{v}$ such that:
@@
The final explanation subgraph is $S_{v}=S^{+}_{v}$.
Mitigating Hallucination.
LLM-generated explanations may hallucinate, especially when multiple node embeddings are projected within a single prompt. We address this issue through two complementary mechanisms within our framework:
  • Prompt templating with constrained node sets: To reduce hallucinations, we partition $\mathcal{T}^{\phi}_{v}$ into fixed-size batches and sequentially prompt them to the LLM, instructing it to generate explanations that reference only the nodes in each batch.
  • Post hoc filtering: After the LLM generates the explanations, any hallucinated nodes—the nodes absent in $\mathcal{T}^{\phi}_{v}$—are filtered out in a post-processing step. This filtering ensures that the final explanation remains relevant and grounded in the actual GNN model’s predictions, as detailed in Section 4.3.
4.4 Faithfulness to Gnn Predictions
Beyond constraining hallucinations, we also aim to ensure that the generated explanations remain faithful to the Gnn’s predictions. Accounting for context aligns the GNN embeddings with those of the LLM; however, this alignment alone does not guarantee that the GNN’s decisions are faithful to the LLM’s. We address faithfulness in two ways.
First, we explicitly include the Gnn’s predicted label for the target node in the input prompt (Sec. 4.2), so that the LLM conditions its reasoning on the model’s actual decision. This provides the LLM with a clear natural-language interpretation of the prediction target.
Second, we introduce Gspell+, which directly optimizes soft prompts to ensure faithfulness to the Gnn’s predictions. Assume the Gnn is a predictive model $\Phi=h\circ g$ where $g:G\rightarrow\mathcal{E}$ is a message-passing encoder and $h:\mathcal{E}\rightarrow Y$ is a decoder s.t. $\mathcal{E}\in\mathbb{R}^{m}$ and $Y\in\mathbb{R}^{C}$ are the latent and output spaces, and $C$ is the number of classes. If $\Phi$ is well optimized, the conditional entropy $H(Y\mid\mathcal{E})$ is low due to the pretraining stage, which is equivalent to increasing the mutual information $I(Y;\mathcal{E})$. Therefore, a good projector $\Pi$ has to not only align the Gnn embeddings with the LLM embeddings, but also have the projections share high mutual information with the output of the Gnn.
Along with the previous alignment objectives $\mathcal{L}_{\text{context}}$ and $\mathcal{L}_{\text{contrast}}$, we further consider increasing the mutual information $I(\hat{y}_{v};Z_{v})$, where $\hat{y}_{v}=h(g(v))\in\mathbb{R}^{C}$ by minimizing the InfoNCE loss van den Oord et al. (2019). Since $Z_{v}$ and $\hat{y}_{v}$ do not live in the same space, we use another projector $\Pi^{\prime}:\mathbb{R}^{k\times h}\rightarrow\mathbb{R}^{C}$ to back-propagate through the original projector $\Pi:\mathbb{R}^{m}\to\mathbb{R}^{k\times h}$ the faithfulness to the prediction. Therefore, writing $Z^{\prime}_{v}=\Pi^{\prime}(Z_{v})$, the final objective for Gspell+ becomes:
$\mathcal{L}=\beta_{1}\;\mathcal{L}_{\text{context}}+\beta_{2}\mathcal{L}_{\text{contrast}}+\beta_{3}\mathcal{L}_{\text{InfoNCE}}\text{ where}$
$\mathcal{L}_{\text{InfoNCE}}=$
$-\mathbb{E}\left[\log\frac{\exp\left(\frac{\text{cos}(Z^{\prime}_{v},\hat{y}_{v})}{\tau}\right)}{\exp\left(\frac{\text{cos}(Z^{\prime}_{v},\hat{y}_{v})}{\tau}\right)+\underset{u\ \neq v}{\sum}\exp\left(\frac{\text{cos}(Z^{\prime}_{v},\hat{y}_{u})}{\tau}\right)}\right]$
Amazon Cora Liar WikiCS
Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size
GNNExplainer 94.5 $\pm$ 1.5 5.71 $\pm$ 0.01 96.5 $\pm$ 0.6 17.91 $\pm$ 0.01 100.0 $\pm$ 0.00 474.01 $\pm$ 0.00 99.0 $\pm$ 0.0 1930.89 $\pm$ 0.0
PGExplainer 73.8 $\pm$ 5.6 4.52 $\pm$ 0.13 82.0 $\pm$ 1.0 8.96 $\pm$ 0.40 100.0 $\pm$ 0.0 *1.0 90.6 $\pm$ 0.0 295.89 $\pm$ 0.0
TAGExplainer 65.1 $\pm$ 4.4 1.03 $\pm$ 0.00 78.5 $\pm$ 0.9 1.06 $\pm$ 0.00 100.0 $\pm$ 0.0 1.31 $\pm$ 0.00 76.4 $\pm$ 1.1 1.51 $\pm$ 0.00
@@ 4.4 Faithfulness to Gnn Predictions
Random 91.0 $\pm$ 1.1 7.28 $\pm$ 0.02 93.7 $\pm$ 0.5 18.67 $\pm$ 0.01 100.0 $\pm$ 0.0 469.36 $\pm$ 0.03 93.3 $\pm$ 0.0 994.76 $\pm$ 0.0
LLM 86.0 $\pm$ 1.2 4.90 $\pm$ 0.23 94.5 $\pm$ 1.0 3.27 $\pm$ 0.31 100.0 $\pm$ 0.0 2.22 $\pm$ 0.04 91.4 $\pm$ 1.0 21.19 $\pm$ 0.30
Gspell+ 92.0 $\pm$ 1.8 4.24 $\pm$ 0.50 85.8 $\pm$ 1.8 2.10 $\pm$ 0.11 100.0 $\pm$ 0.0 1.10 $\pm$ 0.08 86.3 $\pm$ 1.7 12.44 $\pm$ 0.27
Gspell 91.1 $\pm$ 2.3 4.25 $\pm$ 0.54 86.5 $\pm$ 0.0 2.07 $\pm$ 0.07 100.0 $\pm$ 0.0 1.11 $\pm$ 0.08 86.5 $\pm$ 0.0 12.80 $\pm$ 0.0
@@
LLM Prompt You are analyzing Amazon product reviews and their co-purchase neighborhoods to understand why it has been classified under a specific category. Your task is to determine which of the neighboring products does (not) support the classification of the target product. Target Product ID: 160 Predicted Category: Sports & Outdoors
Product Embedding Representation:
<<BEGIN TARGET KEYWORDS>> PROJECTED EMBEDDING <<END TARGET KEYWORDS>>
Neighboring Products:
Each Product below is described by keywords.
Product 3: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 4: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 5: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 10: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
Product 12: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
LLM Response
P3 Summary: This product seems to be a fitness accessory, possibly a yoga mat or a resistance band, given its mention of “workout” and “stretching”. YES
P4 Summary: This product appears to be a nutrition supplement, possibly a protein powder or a vitamin, given its mention of “muscle” and “energy”. NO
P5 Summary: This product seems to be a fitness tracker or a smartwatch, given its mention of “fitness” and “activity tracking”. YES
P10 Summary: This product seems to be a health and wellness product, possibly a meditation app or a mindfulness journal, given its mention of “relax” and “meditate”. NO
P12 Summary: This product appears to be a sports accessory, possibly a water bottle or a towel, given its mention of “hydrate” and “dry off”. YES
Instructions For each neighboring product: 1. Summarize the similarity between supporting products in one sentence. 2. Clearly state whether this product supports the classification of the Target Product into category Sports & Outdoors. Format:
Product <ID>:
Summary: <One sentence>.
Support: YES or NO — Does this product support classification into Sports & Outdoors?
Base reasoning only on the keywords and proximity to the target product.
Figure 2: Left: prompt with category and embeddings highlighted. Right: full model response with summaries in the middle and YES/NO verdicts aligned on the right.
@@ 5 Experiments
We demonstrate the effectiveness of Gspell in various settings. Our code is publicly available at: https://github.com/pbaghershahi/GSPELL.git.
@@ 5.1 Experimental setup
Evaluation Metrics. Having an explainer $\Psi$ of GNN $\Phi$, a dataset $\mathcal{G}$ of $\mathcal{V}$ nodes, and denoting explanation subgraph by $S_{v}=\Psi(v,G)$, we use the following metrics in the experiments:
1) Fidelity:
Fidelity captures how well an explainable model reproduces the GNN model’s logic in its predictions. Mathematically, we have:
$Fid=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in V}\mathbb{I}[\Phi(v,S_{v})=\Phi(v,G)]$
where $\mathbb{I}[\cdot]$ is the indicator function.
2) Size:
Smaller explanation sizes are easier to interpret. The explanation size measures the compactness of the explanation subgraph, i.e., the number of nodes in the explanation, as:
$Size=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in\mathcal{V}}|S_{v}|$
For a more comprehensive evaluation, we adopt the perturbation-based framework of recent natural language explanation literature Atanasova et al. (2023); Siegel et al. (2024).
3) Comprehensiveness:
Comprehensiveness measures the drop in prediction accuracy after removing the explanation subgraph. Let $G\setminus S_{v}$ denote the graph obtained by removing the nodes in $S_{v}$. Then, we define comprehensiveness as:
$Comp.=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},\,v\in V}\mathbb{I}\!\left[\Phi(x_{v},G)=y_{v}\right]$-
$\mathbb{I}\!\left[\Phi\!\left(x_{v},G\setminus S_{v}\right)=y_{v}\right],$
4) Alignment:
Alignment evaluates whether the nodes identified as opposing the prediction, the set $S_{v}^{-}$, are indeed unsupportive. We define alignment as the drop in prediction accuracy after removing these nodes:
$Align.=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},\,v\in V}\mathbb{I}\!\left[\Phi(v,G)=y_{v}\right]$-
$\mathbb{I}\!\left[\Phi\!\left(v,G\setminus S_{v}^{-}\right)=y_{v}\right].$
5) Random Baseline:
The random baseline measures the drop in prediction accuracy after removing a random node set, the same size as the explanation subgraph. Let $S_{v}^{\mathrm{rand}}\subseteq T_{v}^{\phi}$ be a uniformly sampled subset such that $|S_{v}^{\mathrm{rand}}|=|S_{v}|$, we define:
$Rand.=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},\,v\in V}\mathbb{I}\!\left[\Phi(v,G)=y_{v}\right]$-
$\mathbb{I}\!\left[\Phi\!\left(v,G\setminus S_{v}^{\mathrm{rand}}\right)=y_{v}\right].$
Overall, higher values are preferred for Fidelity and Comprehensiveness, but lower values are preferred for Size and Alignment. The Random Baseline serves as a reference, where we expect Comprehensiveness to be significantly higher than the Random Baseline.
@@
Datasets. We use real-world TAG datasets to evaluate the performance of our method: Cora Sen et al. (2008), WikiCS Mernyei and Cangea (2022), Amazon-Product Feng et al. (2024), and Liar Wang (2017). Additional details of the datasets are provided in Appendix B.
Models. We consider a GCN Kipf and Welling (2017b) as the base GNN model for the node classification task, and the LLM model is Meta’s Llama-3.1-8B-Instruct, which remains frozen in our experiments. Full experimental details are provided in Appendix D. We provide experiments with other LLM models in Appendix C.2 and for different GNN architectures in Appendix C.1.
Dimension Cora Amazon-Product
$M_{1}\uparrow$ $M_{2}\uparrow$ $\Delta$ $M_{1}\uparrow$ $M_{2}\uparrow$ $\Delta$
Understandability $2.9_{\scriptstyle\pm 1.23}$ $3.2_{\scriptstyle\pm 1.45}$ $0.3$ $3.33_{\scriptstyle\pm 1.46}$ $3.25_{\scriptstyle\pm 1.33}$ 0.08-0.08
Trustworthiness $2.7_{\scriptstyle\pm 1.10}$ $3.1_{\scriptstyle\pm 1.06}$ $0.4$ $3.30_{\scriptstyle\pm 1.56}$ $3.05_{\scriptstyle\pm 1.41}$ 0.25-0.25
Insightfulness $3.3_{\scriptstyle\pm 0.74}$ $3.1_{\scriptstyle\pm 0.95}$ 0.2-0.2 $2.95_{\scriptstyle\pm 1.34}$ $3.25_{\scriptstyle\pm 1.17}$ $0.30$
Satisfaction $2.2_{\scriptstyle\pm 1.16}$ $2.7_{\scriptstyle\pm 0.92}$ $0.5$ $3.08_{\scriptstyle\pm 1.46}$ $3.05_{\scriptstyle\pm 1.36}$ 0.03-0.03
Confidence $0_{\scriptstyle\pm 1.49} 3.3$$3_{\scriptstyle\pm 1.49}$ $0.3$ $3.13_{\scriptstyle\pm 1.47}$ $3.03_{\scriptstyle\pm 1.33}$ 0.10-0.10
Convincingness $3.3_{\scriptstyle\pm 1.57}$ $3.7_{\scriptstyle\pm 1.64}$ $0.4$ $3.08_{\scriptstyle\pm 1.46}$ $2.93_{\scriptstyle\pm 1.49}$ 0.15-0.15
Communicability $3.0_{\scriptstyle\pm 1.60}$ $3.1_{\scriptstyle\pm 1.41}$ $0.1$ $2.88_{\scriptstyle\pm 1.28}$ $2.93_{\scriptstyle\pm 1.23}$ $0.05$
Usability $3.3_{\scriptstyle\pm 1.43}$ $3.4_{\scriptstyle\pm 1.42}$ $0.1$ $2.75_{\scriptstyle\pm 1.39}$ $2.80_{\scriptstyle\pm 1.22}$ $0.05$
@@
Table 2: Expert qualitative evaluation of the explanations from Gspell and GNNExplainer on Cora and Amazon-Product. $M_{1}$ denotes the mean scores for GNNExplainer and $M_{2}$ the mean scores for Gspell.
Amazon Cora WikiCS
F (%) $\uparrow$ C (%) $\uparrow$ R (%) A (%) $\downarrow$ S $\downarrow$ F (%) $\uparrow$ C (%) $\uparrow$ R (%) A (%) $\downarrow$ S $\downarrow$ F (%) $\uparrow$ C (%) $\uparrow$ R (%) A (%) $\downarrow$ S $\downarrow$
GNNExplainer 94.5 14.7 6.9 - 5.71 96.5 15.8 7.8 - 17.91 99.0 13.1 13.1 - 1930.89
Gspell 91.1 8.2 4.8 3.0 4.25 86.5 5.8 1.2 7.8 2.07 86.5 4.20 0.0 4.7 12.80
Gspell+ 92.0 8.3 3.5 5.0 4.24 85.8 6.2 1.5 8.2 2.10 86.3 3.8 0.2 5.8 12.44
Table 3: Perturbation-based evaluation of faithfulness. Gspell shows a superior trade-off between faithfulness and conciseness, achieved by isolating critical nodes for human-centered explanations in natural language.
Baselines. We consider the following five baselines: (1) Node: This baseline uses only the target node itself as the explanation subgraph. This represents the minimal possible context and serves as a lower bound on explanation size. (2) Random: We introduce a randomized baseline that selects a subgraph around each target node as its explanation subgraph. Specifically, for each node, we sample a subgraph of half the size of its computation tree. The results are averaged over 5 samples. (3-5) Graph Explainers: We compare Gspell with GNNExplainer Ying et al. (2019a), PGExplainer Luo et al. (2020) as subgraph-based explainers, which do not involve any LLM-based component. We also include the recent LLM-based graph explanation method, TAGExplainer Pan et al. (2024).
5.2 Quantitative Results
Table 1 shows that the Node baseline is relatively strong, highlighting the critical role of textual attributes in GNN predictions. While GNNExplainer often reaches the highest fidelity, this metric alone can be misleading. Since GNNs aggregate neighborhood data, high fidelity can be trivially achieved by simply outputting the entire induced subgraph.
In homophilic graphs, where neighbors often share the same category, selecting a broad neighborhood effectively replicates the GNN’s original context. This is evidenced by the Random baseline, which achieves fidelity scores comparable to GNNExplainer. In these settings, attaining high fidelity is trivial if the explanation subgraph is large.
Across all datasets, Gspell achieves a superior fidelity-size trade-off: despite marginally lower fidelity than GNNExplainer and the LLM baseline, it produces significantly smaller subgraphs. Unlike Random or Node baselines, which either inflate size or restrict context, Gspell leverages projected GNN embeddings to distill rich textual semantics into concise, interpretable explanations.
Comparing Gspell+ to Gspell highlights the impact of the InfoNCE objective: while it further reduces explanation size, it occasionally lowers fidelity. The objective’s effectiveness scales with the number of negative samples; for instance, it outperforms on Amazon-Product compared to Cora because a higher class count provides a tighter mutual information bound.
Perturbation-based Evaluation.
To further assess the faithfulness of the generated explanations, we adopt perturbation-based evaluation.
[…]
Unlike baselines that provide binary-mask explanations, Gspell ’s 3-way classification—Support, Oppose, and Neutral—uniquely enables the Alignment metric to differentiate between contradicting evidence and irrelevant noise to detect hallucinations. Across all datasets, Gspell’s Alignment scores are low, particularly on Amazon, confirming successful detection of irrelevant nodes.